# GateLens: A Reasoning-Enhanced LLM Agent for Automotive Software Release Analytics

### v2

@@ IV-D Effectiveness of the RA module (RQ3)
The RA module maintained consistent performance for simpler queries, demonstrating its versatility across varying complexity levels. By transforming natural language into precise, logical representations, the RA module serves as a critical bridge between user intent and code execution. This translation process enables the code generator to produce accurate, efficient executable code for data analysis tasks.
These results establish RA as a fundamental component in our system’s architecture, significantly enhancing its reasoning capabilities and ensuring reliable code generation across diverse query types and complexity levels. This enables the system to deliver accurate and efficient solutions for tabular data analysis tasks.
IV-E Impact of Few-Shot Examples (RQ4)
@@
The results demonstrate that GateLens relies heavily on its RA translation process, achieving optimal performance even in a 0-shot setting without any examples. Interestingly, when a small number of examples are added, GateLens becomes slightly biased toward them, leading to a minor degradation in performance. However, it regains optimal performance with additional examples as the system learns to generalize better. In contrast, CoT+SC’s performance heavily depends on in-context examples, showing suboptimal results without sufficient few-shot examples. While CoT+SC’s performance improves with more examples, this approach is fundamentally inefficient and impractical. Increasing example count expands input context size, leading to higher latency due to the quadratic complexity of Transformer-based models, particularly problematic for real-time applications. This also escalates operational costs through increased token usage (higher cloud API fees) and computational demands. Moreover, the larger context risks exceeding the maximum length, which can lead to truncation or lost information, compromising the model’s response quality. In long contexts, critical content may be ignored, resulting in a phenomenon known as “lost in the middle,” which adversely affects overall performance (Liu et al., 2023).
The results further highlight the critical role of RA in enabling effective query handling. GateLens, through its RA-based translation process, achieves robust performance without requiring extensive in-context learning. GateLens’s independence from examples not only ensures consistent performance across diverse queries but also offers practical benefits by mitigating issues related to context length limitations and accordingly reducing computational and financial resource consumption, enhancing system scalability, and minimizing maintenance needs for adaptation to new tasks. These advantages position GateLens as a robust and efficient solution for automating the analysis of tabular data in dynamic environments.
@@
V Industrial Deployment: Lessons Learned
TABLE V: GateLens (zero-shot) vs CoT+SC (few-shot) performance across different roles on the second benchmark. For each role tested, CoT+SC was trained using examples from the other two roles only (leave-one-role-out approach).
Roles # Queries GateLens CoT+SC (Few-shot with All Roles) CoT+SC (Few-shot with Leave-One-Role-Out)
(244 in total) F1 Score F1 Score Without Mechanic Without Project Without Software
Mechanically-oriented 36 94.25% 80.60% 73.85% $\downarrow$ (-6.75%) 86.57% 80.60%
Project-oriented 193 80.21% 78.93% 78.93% 76.22% $\downarrow$ (-2.71%) 78.40%
Software-oriented 15 100% 100% 100% 100% 82.76% $\downarrow$ (-17.24%)
The deployment of GateLens at a partner automotive company has provided valuable insights into integrating AI-assisted analytics into complex industrial workflows, specifically for streamlining decision-making in automotive software release validation.
Automotive software integration at the company typically occurs across three hierarchical stages: subsystem (control unit), system (multiple control units), and full vehicle levels. Each stage involves extensive testing, with results stored in a central database. Critical Go/No-Go decisions are made at these stages to determine whether a release meets quality thresholds. However, stakeholders from diverse backgrounds—including project managers, mechanical engineers, and software engineers—must query the raw data to evaluate product quality. Many lack expertise in data analytics, creating bottlenecks and delays in the decision-making process.
Previously, these analytics were managed by a small team of 2–3 full-time analysts, who were often overwhelmed by the volume and diversity of requests. Scaling the team to meet the current demand would have required tripling its size. GateLens addresses this challenge by automating much of the workload, enabling more efficient decision-making. Currently, GateLens is in an extended pilot phase, supporting a pool of 60-80 users. The analytics team has transitioned to a support role, helping stakeholders articulate their needs into clear, actionable prompts for the system. User adoption of GateLens has progressed in phases:
  • Small-Scale Pilot: The initial deployment within the analytics team established benchmarks.
  • Expanded Pilot: Five additional users from varied backgrounds contributed to refining the benchmarks.
  • Wider Rollout: The current phase involves a larger group of 60–80 users. Feedback has been highly positive, with stakeholders recognizing GateLens’s ability to simplify and accelerate complex analyses.
Since the launch, the number of both new and recurring users has grown, encompassing diverse roles and types of queries, thereby demonstrating the tool’s increasing utility and trust. GateLens significantly reduces the time and effort required for complex analyses, but the shift towards automation also requires users to take on more responsibility in defining and clarifying their needs. The transition from a primarily supportive tool to a more fully automated system is ongoing, demanding a gradual approach with careful calibration to ensure the tool continues to meet evolving needs.
The diversity in stakeholders’ needs and backgrounds is an inevitable factor, leading to a wide range of requirements and queries when interacting with analytics systems. Our preliminary agentic system based on CoT and SC has performed well when serving a specific group of stakeholders. However, as we opened the system to a broader audience, the variety of queries increased substantially. Systems that rely heavily on few-shot learning face significant challenges in these situations, as they are limited by the relatively small set of few-shot examples, making it difficult to handle a wider range of potentially unexpected inputs. Similarly, fine-tuned models often struggle to adapt to dynamic and diverse environments. Improving generalization is essential to ensure the scalability and reliability of analytics systems in these complex, domain-specific contexts. In our design and development work, we prioritized the system’s ability to generalize effectively and meet the needs of a broader and more diverse group of users.
To explore the system’s generalizability, we categorized the roles within the company into three groups: mechanically-oriented, project-oriented, and software-oriented roles. Mechanically-oriented roles typically focus on truck-specific data filtering. Project-oriented roles often combine meta-queries with conditional filters for release management and statistical analysis. Software roles emphasize truck software applications and user functions. We can see from Table V, both GateLens (zero-shot) and CoT+SC (few-shot) exhibit differences in system performance across these groups, which is likely stemming from the complexity and variety of their typical queries. Nevertheless, the results demonstrate that GateLens is capable of supporting all groups to a high degree. To further evaluate CoT+SC’s dependency on few-shot examples, we conducted a leave-one-role-out experiment. In this approach, examples from a specific role are excluded in each iteration. For instance, ’without software’ indicates that all examples from the software-oriented role have been removed, while the total number of examples is maintained by substituting them with examples from other roles. This highlights the potential challenges with the robustness and generalizability of techniques that rely on few-shot examples. This is a crucial factor to consider in industries where diverse teams collaborate and a wide range of queries may arise.
The impact of automated systems, such as GateLens, on the release process has been substantial. Compared to the previous manual process, GateLens has reduced the time required for Go/No-Go analytics by more than 80%, significantly improving operational efficiency. Stakeholders can now focus on high-level decision-making, freed from the burden of data preparation and analysis.
A key advantage of GateLens lies in its domain-specific design. Unlike general-purpose tools like TaskWeaver (Qiao et al., 2023) or AutoGen (Wu et al., 2023), GateLens is tailored to automotive workflows, making it easier to understand, debug, and adapt to automotive common procedures. This focus on domain relevance ensures that the system aligns more closely with stakeholders’ needs while providing reliable and nuanced support.
In summary, the deployment of GateLens demonstrates how domain-specific AI solutions can transform critical workflows in the automotive sector. By automating labor-intensive processes and enhancing decision-making, GateLens has delivered measurable improvements in efficiency and user satisfaction. However, its success depends on ongoing refinement and careful management of the transition to full(er) automation. Balancing automation with user empowerment remains crucial, particularly in a complex industry like automotive, where diverse stakeholder needs must be met.
GateLens represents a promising step forward, showcasing the potential of AI-driven systems to improve not only the automotive domain but also other industries requiring robust, scalable solutions for intricate processes.
VI Related Work
General-purpose LLMs are primarily designed for and trained on natural languages. Working with tabular data requires specialized adaptations to effectively handle its structured and heterogeneous nature (Fang et al., 2024; Wang et al., 2024b; van Breugel and van der Schaar, 2024). First, the structured tabular data is typically transformed into serialized text. The performance of the LLM may depend on this transformation (Min et al., 2024). Subsequently, the serialized text data is used as input to the LLM for various tasks, such as question-answering, summarization, or logical reasoning. Common approaches to improve LLM performance include prompt engineering, pre-training, fine-tuning, and Retrieval-Augmented Generation (RAG).
Pre-training and fine-tuning (Zhang et al., 2023b; Parthasarathy et al., 2024; VM et al., 2024; Dong et al., 2022; Hegselmann et al., ) often face scalability concerns. Although resource-efficient training techniques have been proposed to mitigate the substantial computational demands of LLMs (Han et al., 2024; Lin et al., 2024), in safety-critical applications with evolving data and requirements, training LLMs presents significant challenges due to the constant need for rigorous validation and verification. This ongoing necessity substantially increases resource demands for development and maintenance, potentially exceeding the capacities of many companies. Techniques such as RAG have been employed to dynamically integrate external knowledge bases during inference, reducing the need for frequent model updates (Zhao et al., 2024; Gao et al., 2023). However, such methods can pose challenges in safety-critical industrial settings as well, since both retrieval modules and model components must undergo synchronized updates to maintain relevance, reliability, and compliance with validation and verification requirements. Costs would also be especially high with fine-tuning since re-tuning would be needed when new and improved base LLMs are released and should be incorporated.
Prompt engineering techniques are among the most resource-efficient methods for improving LLM output (Sahoo et al., 2024; Jin and Lu, 2023). From a user standpoint, when the input is natural language, prompting techniques can be broadly categorized based on the type of language generated by the LLM. These include outputs in natural language, structured languages (Li et al., 2023), or symbolic languages. When the generated language is natural language, LLMs often fail to consistently follow instructions, particularly when the instructions are complex or require precise, step-by-step execution (Pham et al., 2024). This inconsistency arises because natural language, while flexible and expressive, can be ambiguous and prone to misinterpretation by LLMs. Structured languages include general-purpose languages (e.g. Python) (Ye et al., 2024), query languages (e.g. SQL) (Li et al., ; Dong et al., 2023; Mouravieff et al., 2024), configuration formats (e.g. YAML or JSON), or other Domain-Specific Languages (DSLs) (Glenn et al., 2024; Dai et al., 2024). These languages are subsequently interpreted and/or executed by either external tools, the same LLM, or another LLM agent. This approach offers significant advantages, as it enables precise execution of tasks. Another popular type of output is symbolic languages. Literature shows that symbolic representations provide a more rigorous framework for articulating premises and intent, which can enhance reasoning capabilities (Pan et al., 2023).
In this paper, we introduce a novel prompt-only (training-free) approach that bridges natural language and executable code through RA, a symbolic formalism designed for relational modeling and ideally suited for analyzing tabular data. Unlike prior work that often relies on complex multi-agent planning, our approach leverages RA as a lightweight intermediate representation to enable precise query normalization, disambiguation of natural language input, and efficient code generation. RA acts as an abstraction layer that can target multiple execution backends (e.g., Python, SQL), providing adaptability across systems. In GateLens, we generate Python code to support practical industrial deployment and high-performance execution. GateLens is training-free, feed-forward (single-pass, without looping or multi-agent orchestration), and thus easier to verify, trace, maintain, and trust – qualities critical for safety-critical industrial applications.
VII Discussion and Conclusions
This study introduced GateLens, a reasoning-enhanced LLM agent designed for automotive software release validation. By introducing RA as an intermediate representation before code generation, Gatelens addresses the ”Unfaithful Chain-of-Thought (CoT) reasoning” problem in code generation, where reasoning steps in CoT explanations do not accurately reflect the model’s actual thought process (Turpin et al., 2023). Specifically, GateLens divides the analysis into two steps: (1) natural language queries are first translated into RA expressions, and (2) these RA expressions are then converted into executable code. We use Python as our target language in step (2) due to its widespread use in our partner company and the model’s strong performance in Python, which benefits from more extensive training data. However, this step is also compatible with RA-to-SQL generation, enabling flexibility across backends. The inclusion of the RA reasoning module is a critical factor in improving robustness and scalability, as evidenced by superior F1 scores in both benchmarking and industrial evaluations.
In real-world deployment serving 60-80 users, GateLens demonstrates significant practical value through its user-friendly interface and robust query processing capabilities, with users particularly appreciating the flexibility to input, debug, and refine queries easily. This marks a substantial advancement in industrial data interaction, successfully handling complex and ambiguous queries while providing practical support for faster decision-making in safety-critical software release processes. GateLens demonstrates significant practical advancements by reducing analysis time by over 80% while maintaining high accuracy in test result interpretation, impact assessment, and release candidate evaluation.
Our implementation insights highlight the advantages of focusing on training-free and single-pass agent systems by foregrounding the perception phase in the code generation pipeline. This modular architecture opens opportunities for incorporating emerging LLM capabilities while preserving the system’s practical utility in safety-critical industrial applications. A phased deployment program is ongoing and shows that multiple stakeholder groups can be supported, though the evolving roles and analytical needs require continued refinement.
Future work will address current limitations by expanding domain applicability beyond automotive software and testing alternative LLM configurations to enhance reliability across other safety-critical industries. This approach demonstrates the potential for LLM-based solutions to transform industrial data analysis workflows while maintaining the reliability standards required for critical applications.
VIII Threats to Validity
The validity of our findings is subject to several potential threats. First, this framework depends on well-defined, static schemas for accurate query decomposition, which fundamentally limits its effectiveness in environments with incomplete or frequently changing schemas. Second, the generalizability of GateLens beyond the automotive software release domain is limited, as the system relies on domain-specific relational modeling and datasets, necessitating further validation in other safety-critical industries. Third, the benchmarks and query scenarios used for evaluation, derived from historical data and real-world queries, may not fully capture the diversity and complexity of potential use cases, which could impact the robustness of the system in broader deployments. Finally, the system’s performance depends on the specific LLM configurations used, such as GPT-4o and Llama 3.1 70B, and their ability to interpret and generate RA expressions. Future work will address these limitations through broader domain testing, expanded evaluation scenarios, and alternative LLM configurations.
Acknowledgments
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.

### v3

@@ 4.4 Effectiveness of the RA module (RQ3)
The RA module maintained consistent performance for simpler queries, demonstrating its versatility across varying complexity levels. By transforming natural language into precise, logical representations, the RA module serves as a key bridge between user intent and code execution. This translation process enables the code generator to produce accurate, efficient executable code for data analysis tasks.
RQ3 findings: The RA module is a critical component for handling query complexity. Removing it results in substantial performance degradation (over 27% for complex queries), confirming that translating natural language to RA provides necessary structural guidance for accurate code generation.
4.5 Role of Few-Shot Examples (RQ4)
@@
The results demonstrate that GateLens relies heavily on its RA translation process, achieving 100% F1 even in a 0-shot setting without any examples. Interestingly, when a small number of examples are added, GateLens becomes slightly biased toward them, leading to a minor degradation in performance (dropping to 95.92% at 2 examples). However, it regains optimal performance at 3 examples and remains at 100% thereafter. In contrast, CoT+SC’s performance heavily depends on in-context examples, achieving only 42% F1 with 2 examples and showing suboptimal results without sufficient few-shot examples. Performance improved steadily with more examples, reaching 84.57% at 50 examples.
@@
RQ4 findings: GateLens does not require many-shot examples to achieve high performance, delivering optimal results in a 0-shot setting. This contrasts with CoT+SC, which depends on carefully curated examples. This independence from few-shot examples makes GateLens more efficient for real-time applications by reducing context size and computational overhead.
5 Qualitative Assessment Examples
Table 6: An example of an input query and its corresponding relational algebra expressions translated using a domain-specific database schema, along with the executable code generated from the relational algebra and the resulting output produced when executed over the database.
Stage Content
Input Query Compute the relative frequency of not run tests for each RC and function
Relational Step 1: $R0\leftarrow\sigma_{\text{result\_status}=\text{"NotRun"}}(\text{df})$
Algebra Step 2: $R1\leftarrow\gamma_{\text{releasecandidate},\text{function},\text{countNotRun}\rightarrow\text{count}}(R0)$
Step 3: $R2\leftarrow\gamma_{\text{releasecandidate},\text{function},\text{totalcount}\rightarrow\text{count}}(\text{df})$
Step 4: $R3\leftarrow R1\bowtie_{\text{R1.releasecandidate}=\text{R2.releasecandidate}\wedge\text{R1.function}=\text{R2.function}}R2$
Step 5: $\text{Result}\leftarrow\gamma_{\text{R3.releasecandidate},\text{R3.function},\text{relativeFrequency}\rightarrow\frac{\text{R3.count}}{\text{R3.totalcount}}}(R3)$
Python Code